### v1

# A note on the cohomology of moduli spaces of local shtukas

Abstract.
We study localized versions of spectral action of Fargues–Scholze, using methods from higher algebra. As our main motivation and application, we deduce a formula for the cohomology of moduli spaces of local shtukas under certain genericity assumptions, and discuss its relation with the Kottwitz conjecture.
1. Introduction
The purpose of this note is to explicate a formula for the cohomology of moduli spaces of local shtukas that can be derived from the recent work of Fargues–Scholze [FS21], and to note some consequences. Let $p\neq\ell$ be distinct primes. The theory of local Shimura varieties, which began with examples such as the Lubin–Tate and Drinfeld towers, and continued with the moduli spaces of $p$-divisible groups of Rapoport–Zink [RZ96], has reached a stage of maturity exceeding that of the global theory of Shimura varieties with Scholze’s general definition of moduli spaces of local shtukas [SW20]. For any connected reductive group $G/\mathbb{Q}_{p}$, any element $b$ in Kottwitz’s set $B(G)$, and any finite collection $\mu_{\bullet}=(\mu_{i})_{i\in I}$ of conjugacy classes of cocharacters $\mu_{i}$ of $G$, Scholze defines a tower
$(\operatorname{Sht}_{(G,b,\mu_{\bullet}),K})_{K}$
of diamonds, with inverse limit $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$, where $K\subseteq G(\mathbb{Q}_{p})$ runs through the compact open subgroups. When $\mu_{\bullet}=\mu$ is a single minuscule cocharacter, the $\operatorname{Sht}_{(G,b,\mu),K}$ are smooth rigid spaces referred to as local Shimura varieties. The space $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$ carries commuting actions of $G(\mathbb{Q}_{p})$ and $G_{b}(\mathbb{Q}_{p})$, where $G_{b}$ is the inner form of a Levi subgroup of the quasisplit form of $G$ canonically attached to the datum $(G,b)$. For any irreducible admissible $G_{b}(\mathbb{Q}_{p})$-representation $\rho$ over $\overline{\mathbb{Q}}_{\ell}$, one may define the “$\rho$-isotypic part of the $\overline{\mathbb{Q}}_{\ell}$-intersection cohomology of $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$”, which we will denote by
$R\Gamma(G,b,\mu_{\bullet})[\rho].$
We note that this definition naturally incorporates a shift making 00 the ‘middle degree’. It carries commuting actions of $G(\mathbb{Q}_{p})$ and $W_{E_{\bullet}}:=\prod_{i\in I}W_{E_{i}}$ (where $W_{E_{i}}$ is the Weil group of the reflex field $E_{i}$ of $\mu_{i}$) and is a bounded complex of finite length admissible $G(\mathbb{Q}_{p})$-representations. Roughly speaking, the local Langlands conjecture associates an $L$-parameter $\phi=\phi_{\rho}:W_{\mathbb{Q}_{p}}\to(\widehat{G}\rtimes W_{\mathbb{Q}_{p}})(\overline{\mathbb{Q}}_{\ell})$ with $\rho$, as well as an $L$-packet $\Pi_{\phi}(G)$ of irreducible admissible representations of $G(\mathbb{Q}_{p})$. Furthermore, there should be a relation between $\operatorname{Irr}(S_{\phi},\chi_{b})$ and $\Pi_{\phi}(G)$. Here $S_{\phi}$ is the centralizer in $\widehat{G}$ of the image of $\phi$, and $\operatorname{Irr}(S_{\phi},\chi_{b})$ denotes the set of irreducible representations of $S_{\phi}$ on which $Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}\subseteq S_{\phi}$ acts by a certain character $\chi_{b}$ determined by $b$. Let $V$ be the dual of the irreducible representation of $\prod_{i\in I}\widehat{G}\rtimes W_{E_{i}}$ with extreme weight $\mu_{\bullet}$. The underlying vector space of $V$ also carries an action of $W_{E_{\bullet}}$ coming from $\phi$, and this action makes it into a $S_{\phi}\times W_{E_{\bullet}}$-representation that we will denote by $V_{\phi}$. Define
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho):=\sum_{n}(-1)^{n}H^{n}(R\Gamma(G,b,\mu_{\bullet})[\rho])$
in the Grothendieck group of $G(\mathbb{Q}_{p})\times W_{E_{\bullet}}$-representations. The following is then the natural generalization of the Kottwitz conjecture (see e.g. [RV14, §7.1]), together with a folklore vanishing conjecture.
Conjecture 1.1.
Assume that $b$ is basic and $\phi=\phi_{\rho}$ is elliptic, i.e. $\phi$ is semisimple and $S_{\phi}/Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}$ is finite. In this case the local Langlands correspondence predicts a surjection $\operatorname{Irr}(S_{\phi},\chi_{b})\to\Pi_{\phi}(G)$, which we denote by $\delta\mapsto\pi_{\delta}$.
  1. (1)
    (Kottwitz conjecture) We have $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})}\pi_{\delta}\boxtimes\operatorname{Hom}_{S_{\phi}}(\delta,V_{\phi})$.
  2. (2)
    (Vanishing conjecture) $H^{n}(R\Gamma(G,b,\mu_{\bullet})[\rho])=0$ for $n\neq 0$.
We will refer to the conjunction of parts (1) and (2) as the strong Kottwitz conjecture. When $b$ is not basic, an extension of the Harris–Viehmann conjecture [RV14, Conjecture 8.4] gives a formula for $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)$ in terms of the basic case for smaller reductive groups; we will not elaborate on this further. Our main goal is a formula for $R\Gamma(G,b,\mu_{\bullet})[\rho]$. We recall that [FS21] associates a semisimple $L$-parameter with any irreducibe admissible representation of a connected reductive group; we call this the Fargues–Scholze parameter. It is expected to be semisimplication of the $L$-parameter appearing in the local Langlands conjecture. Our main result is then the following:
Theorem 1.2.
Assume that the Fargues–Scholze parameter $\varphi$ attached to $\rho$ is generous. Then
$R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}C_{\delta}\boxtimes\operatorname{Hom}_{S_{\varphi}}(\delta,V_{\varphi}),$
in the derived category of $G(\mathbb{Q}_{p})\times W_{E_{\bullet}}$-representations. If $\varphi$ is elliptic, then $C_{\delta}$ is a nonzero split bounded complex of finite direct sums of supercuspidal representations of $G(\mathbb{Q}_{p})$ with Fargues–Scholze parameter $\varphi$.
The precise version, which notably includes a formula for each $C_{\delta}$ in terms of $\rho$ and $\delta$, is given in Theorem 3.5 and Corollary 3.8 (and works in positive characteristic as well). A parameter $\varphi$ is called generous if it is semisimple, if no other $L$-parameter has semisimplification $\varphi$, and if $\varphi$ satisfies an additional technical moduli-theoretic condition (see Definition 3.1). Generous parameters are generic on the coarse moduli space and includes the elliptic parameters. Thus, under the expectation that $\varphi=\phi$, the action of $W_{E_{\bullet}}$ on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ is as predicted from the Kottwitz conjecture. We note that our formula may be seen as a (more general) local analogue of [LZ19, Prop. 1.2].
The proof of Theorem 1.2 is given in sections 2 and 3. The main idea is to apply the machinery of higher algebra to the spectral action of Fargues–Scholze, to obtain a version of the spectral action which only sees one $L$-parameter at a time. The general version of this idea is described in §2, and in §3 we use to derive Theorem 1.2. The key point that we wish to make, and which is needed to carry out the proof, is that even though the machinery that we use (monoidal $\infty$-categories and their modules) is highly abstract, it allows you to make computations. This would fail if we tried to work with triangulated categories instead of their $\infty$-categorical enhancements.
Section 4 then gives some applications of Theorem 1.2 to both parts of Conjecture 1.1; we highlighting one such application. When disregarding the $W_{E_{\bullet}}$-action, Conjecture 1.1 was recently proven in [HKW22] under the assumption of a precise form of the local Langlands correspondence. Combining this with Theorem 1.2, one gets the following result.
Theorem 1.3.
Assume the refined local Langlands correspondence [Kal16, Conjecture G], and let $\phi$ be the $L$-parameter attached to $\rho$. Assume further that the Fargues–Scholze parameter $\varphi$ attached to $\rho$ is elliptic, that all $\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})$ are one-dimensional, and that all representations in $\Pi_{\phi}(G)$ are supercuspidal. Then there exists a surjection $\delta\mapsto\pi_{\delta}^{\prime}$ from $\operatorname{Irr}(S_{\varphi},\chi_{b})$ to $\Pi_{\phi}(G)$ such that
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}\pi_{\delta}^{\prime}\boxtimes V_{\varphi,i}$
in the Grothendieck group of $G(\mathbb{Q}_{p})\times W_{E_{\bullet}}$-representations.
In particular, this is the Kottwitz conjecture, up to the comparison of the local Langlands correspondence and the Fargues–Scholze construction.
Acknowledgements
C.J. was supported by Vetenskapsrådet Grant 2020-05016, Geometric structures in the p-adic Langlands program. D.H. was supported by a startup grant through the National University of Singapore.
2. Spectral action with supports
In this section we derive a version of the spectral action, taking supports in the coarse moduli space of $L$-parameters into account. A very special case of this is considered in [FS21], when the support is a connected component of the coarse moduli space. The general statement turns out to be rather formal, using the machinery of higher algebra. We will use the same notation as in [FS21] as much as possible, except that we will work over $\overline{\mathbb{Q}}_{\ell}$ as opposed to the more general $\mathbb{Z}_{\ell}$-algebras $\Lambda$ considered in [FS21] [^1] Working over $\overline{\mathbb{Q}}_{\ell}$, as opposed to over a general field extension of $\mathbb{Q}_{\ell}(\sqrt{q})$, is just a matter of convenience. We would hope that much of our picture carries over to all cases considered in [FS21], but there are additional technical challenges in mixed characteristic..
We start with a quick recap of some of the main players. In what follows, $\ell\neq p$ will be two distinct primes and $G$ be a connected reductive group over a local field $E$ of residue characteristic $p$. The size of the residue field of $E$ will be denoted by $q$. The dual group of $G$ over $\overline{\mathbb{Q}}_{\ell}$ will be denoted by $\widehat{G}$. $\widehat{G}$ carries an action of the Weil group $W_{E}$, which factors through a finite quotient $Q$, and one can form the semidirect product $\widehat{G}\rtimes Q$. In [FS21, §III], the Artin v-stack $\mathrm{Bun}_{G}$ of $G$-bundles on the Fargues–Fontaine curve is defined. Its underlying topological space $|\mathrm{Bun}_{G}|$ is naturally identified with Kottwitz’s set $B(G)$, and for any $b\in B(G)$ we have an immersion
$i^{b}:\mathrm{Bun}_{G}^{b}\to\mathrm{Bun}_{G}.$
The main player on the geometric side of the geometrization of the local Langlands correspondence is the stable $\infty$-category $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\overline{\mathbb{Q}}_{\ell})$ defined in [FS21, §VII.7] and its counterparts $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{Q}}_{\ell})$ on the strata $\mathrm{Bun}_{G}^{b}$, which are equivalent to the derived categories $\mathcal{D}(G_{b}(E),\overline{\mathbb{Q}}_{\ell})$ of smooth representations of $G_{b}(E)$ over $\overline{\mathbb{Q}}_{\ell}$.
Convention 2.1.
Since we will only work over $\overline{\mathbb{Q}}_{\ell}$, we drop it from the notation for the objects on the geometric side, writing $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}):=\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\overline{\mathbb{Q}}_{\ell})$, $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}^{b}_{G}):=\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}^{b}_{G},\overline{\mathbb{Q}}_{\ell})$, $\mathcal{D}(G_{b}(E)):=\mathcal{D}(G_{b}(E),\overline{\mathbb{Q}}_{\ell})$, etc.
For any stable $\infty$-category $\mathcal{D}(\dots)$, its homotopy category will be denoted by $D(\dots)$, and in any ($\infty$-) category $\mathcal{C}$, we will let $\mathcal{C}^{\omega}$ denote the full subcategory of compact objects. $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ carries the action of Hecke operators [FS21, Thm. IX.0.1]: For every finite set $I$ and any algebraic representation $V$ of $(\widehat{G}\rtimes Q)^{I}$ over $\overline{\mathbb{Q}}_{\ell}$, there is an exact functor
$T_{V}:\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})\to\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})^{BW_{E}^{I}}$
which preserves compact objects, where $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})^{BW_{E}^{I}}$ denotes the category of $W_{E}^{I}$-equivariant objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ (see [FS21, §IX.1] for precise definitions). The $T_{V}$ fit together into exact $\operatorname{Rep}(Q^{I})$-linear monoidal functors
$\operatorname{Rep}((\widehat{G}\rtimes Q)^{I})\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})^{\omega})^{BW_{E}^{I}},$
which are functorial as $I$ varies.
On the spectral side, we have the stack of $L$-parameters $Z^{1}(W_{E},\widehat{G})/\widehat{G}$ over $\overline{\mathbb{Q}}_{\ell}$, defined in [FS21, §VIII]. Since it will figure extensively in our discussion and formulas, we simplify the notation by defining
$\mathfrak{X}:=Z^{1}(W_{E},\widehat{G})/\widehat{G}.$
We also let $X^{\Box}:=Z^{1}(W_{E},\widehat{G})$ be the representation variety and let $X:=Z^{1}(W_{E},\widehat{G})\sslash\widehat{G}$ be its GIT quotient (the character variety), which is a good moduli space for $\mathfrak{X}$. The spectral Bernstein center $\mathcal{Z}^{\mathrm{spec}}(G):=\mathcal{O}(X^{\Box})^{\widehat{G}}$ is then the ring of global functions on $\mathfrak{X}$, and we let $\operatorname{Perf}(\mathfrak{X})$ denote the stable $\infty$-category of perfect complexes on $\mathfrak{X}$. By [FS21, Thm. X.0.1], the action of the Hecke operators on $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ is equivalent to an exact $\overline{\mathbb{Q}}_{\ell}$-linear monoidal functor
$\operatorname{Perf}(\mathfrak{X})\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})^{\omega}),$
called the spectral action. Here and below $\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}$ denotes the $\infty$-category of exact $\overline{\mathbb{Q}}_{\ell}$-linear endofunctors, and $\operatorname{End}^{L}_{\overline{\mathbb{Q}}_{\ell}}\subseteq\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}$ denotes the full subcategory of colimit-preserving endofunctors.
Remark 2.2.
[…]
$\operatorname{Perf}(\mathfrak{X})\to\operatorname{Perf}(\mathfrak{X}^{P})\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}^{P}(\mathrm{Bun}_{G},)^{\omega}).$
In particular, they extend uniquely to colimit-preserving actions
$\operatorname{QCoh}(\mathfrak{X})\to\operatorname{QCoh}(\mathfrak{X}^{P})\to\operatorname{End}^{L}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}^{P}(\mathrm{Bun}_{G})),$
and this induces an action $\operatorname{QCoh}(\mathfrak{X})\to\operatorname{End}^{L}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}))$. See [FS21, §IX.5] for precise definitions of the objects above. If $V\in\operatorname{QCoh}(\mathfrak{X})$, we will write $V\ast-$ or $\operatorname{Act}_{V}(-)$ for the corresponding endofunctor on $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$, depending on which notation fits best with the situation at hand.
The spectral action makes $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ into a module for $\operatorname{QCoh}(\mathfrak{X})$ in the sense of higher algebra (we refer to [Lur] for the notions of higher algebra). This module structure will allow us to apply the constructions of higher algebra to $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$. Before doing so, we recall a few more properties. First, the Hecke action is recovered from the spectral action as the composition
$\operatorname{Rep}((\widehat{G}\rtimes Q)^{I})\to\operatorname{Perf}(\mathfrak{X})^{BW_{E}^{I}}\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}))^{\omega})^{BW_{E}^{I}},$
where the second functor is induced from the spectral action. The first functor sends $V\in\operatorname{Rep}((\widehat{G}\rtimes Q)^{I})$ to the vector bundle $V\otimes\mathcal{O}_{X^{\Box}}$ on $X^{\Box}$, with $W_{E}^{I}$-action given by
$W_{E}^{I}\to(\widehat{G}\rtimes Q)^{I}(\mathcal{O}_{X^{\Box}})\to\operatorname{GL}(V\otimes\mathcal{O}_{X^{\Box}}),$
where the first map is the universal homomorphism in each factor, and the $\widehat{G}$-descent datum comes from the diagonal embedding $\widehat{G}\to(\widehat{G}\rtimes Q)^{I}$. By looking at centres, the spectral action induces a ring homomorphism
$\mathcal{Z}^{\mathrm{spec}}(G)\to\mathcal{Z}^{\mathrm{geom}}_{\mathrm{Hecke}}(G)\subseteq\mathcal{Z}^{\mathrm{geom}}(G)$
to the geometric Bernstein center $\mathcal{Z}^{\mathrm{geom}}(G)$ of $G$, landing inside the subring $\mathcal{Z}^{\mathrm{geom}}_{\mathrm{Hecke}}(G)$ of endomorphisms equivariant for the Hecke action [FS21, Def. IX.0.2]. For any object $A\in\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ we have a ring homomorphism $F_{A}:\mathcal{Z}^{\mathrm{spec}}(G)\to\operatorname{End}_{D_{\mathrm{lis}}(\mathrm{Bun}_{G})}(A)$. When $A$ is Schur (i.e. $\operatorname{End}_{D_{\mathrm{lis}}(\mathrm{Bun}_{G})}(A)=\overline{\mathbb{Q}}_{\ell}$), the kernel of $F_{A}$ corresponds to a $\overline{\mathbb{Q}}_{\ell}$-point of $X$, which we refer to as the Fargues–Scholze parameter of $A$. By [Zou22, Theorem 5.2.1], this construction is compatible with that given in [FS21]. For $b\in B(G)$ and $\rho$ an irreducible admissible representation of $G_{b}(E)$, we will write “the Fargues–Scholze parameter of $\rho$” to mean the Fargues–Scholze parameter of $i^{b}_{\ast}\rho$ [^2] We note that using $i_{\natural}^{b}\rho$ instead of $i^{b}_{\ast}\rho$, as is done in [FS21, IX.7], produces the same result (e.g. by the argument in the proof of Proposition 2.6). Moreover, the Fargues–Scholze parameter of $i^{b}_{\ast}\rho$ is the composite of the “correct” Fargues–Scholze parameter of $\rho$ (defined using $i^{1}_{!}\rho$ in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G_{b}})$ as in [FS21, Def. IX.7.1]) and the twisted $L$-embedding $\widehat{G}_{b}\rtimes Q\to\widehat{G}\rtimes Q$, by [FS21, Thm IX.7.2]..
Now consider a derived stack $\mathfrak{Y}$ with a map $f:\mathfrak{Y}\to\mathfrak{X}$, which induces a monoidal functor $f^{\ast}:\operatorname{QCoh}(\mathfrak{X})\to\operatorname{QCoh}(\mathfrak{Y})$. Using the spectral action and $f^{\ast}$, we can then form the tensor product
$\operatorname{QCoh}(\mathfrak{Y})\otimes_{\operatorname{QCoh}(\mathfrak{X})}D_{\mathrm{lis}}(\mathrm{Bun}_{G})$
in higher algebra, which is a $\operatorname{QCoh}(\mathfrak{Y})$-module. The basic idea of this paper is that there should be plenty of interesting objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ which have $\operatorname{QCoh}(\mathfrak{Y})$-structures (for suitable $\mathfrak{Y}$), and that this can be used for concrete computations.
To keep the definitions of this paper close to its theorems, we make our definitions in a restricted setting, where things are more concrete. Consider the character variety $X$ from above. Let $Y$ be a closed subscheme of $X$, and consider its derived pullback $\mathfrak{Y}$ to $\mathfrak{X}$. The structure sheaf $\mathcal{O}_{\mathfrak{Y}}$ is a commutative algebra object of $\operatorname{QCoh}(\mathfrak{X})$. Similarly, the structure sheaf $\mathcal{O}_{Y}$ of $Y$ is a commutative algebra object of $\operatorname{QCoh}(X)$, and pullback gives a natural monoidal functor $\operatorname{QCoh}(X)\to\operatorname{QCoh}(\mathfrak{X})$ sending $\mathcal{O}_{Y}$ to $\mathcal{O}_{\mathfrak{Y}}$. In particular, we may and will also view $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ as module for $\operatorname{QCoh}(X)$.
Definition 2.3.
Let $Y\subseteq X$ be a closed subscheme and let $\mathfrak{Y}$ be its derived pullback to $\mathfrak{X}$.
  1. (1)
    We define $\operatorname{QCoh}^{Y}(\mathfrak{X})$ to be the $\infty$-category $\operatorname{Mod}_{\mathcal{O}_{\mathfrak{Y}}}(\operatorname{QCoh}(\mathfrak{X}))$ of $\mathcal{O}_{\mathfrak{Y}}$-module objects in $\operatorname{QCoh}(\mathfrak{X})$. Equivalently, $\operatorname{QCoh}^{Y}(\mathfrak{X})$ is $\operatorname{Mod}_{\mathcal{O}_{Y}}(\operatorname{QCoh}(\mathfrak{X}))$, regarding $\operatorname{QCoh}(\mathfrak{X})$ as a module for $\operatorname{QCoh}(X)$. We also set $\operatorname{QCoh}^{Y}(X):=\operatorname{Mod}_{\mathcal{O}_{Y}}(\operatorname{QCoh}(X))$.
  2. (2)
    We define $\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}_{G})$ to be the $\infty$-category $\operatorname{Mod}_{\mathcal{O}_{\mathfrak{Y}}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}))$ of $\mathcal{O}_{\mathfrak{Y}}$-module objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$. Equivalently, $\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}_{G})$ is $\operatorname{Mod}_{\mathcal{O}_{Y}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}))$, regarding $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ as a module for $\operatorname{QCoh}(X)$.
When $Y=\{\varphi\}$ is a closed point, we will simply write $\operatorname{QCoh}^{\varphi}(\mathfrak{X})$ and $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$.
We refer to [Lur, §4] for the theory of algebra objects and their modules in higher algebra. By [Lur, Prop. 7.1.1.4], all $\infty$-categories defined in Definition 2.3 are stable $\infty$-categories (using exactness of the spectral action). Moreover, $\operatorname{QCoh}^{Y}(\mathfrak{X})$ is a symmetric monoidal $\infty$-category and $\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}_{G})$ is a module for it. Both may also viewed as modules for $\operatorname{QCoh}^{Y}(X)$ via $\operatorname{QCoh}^{Y}(X)\to\operatorname{QCoh}^{Y}(\mathfrak{X})$. We then have the following well known fact, which is a standard application of Lurie’s Barr–Beck Theorem [Lur, Thm. 4.7.3.5].
Proposition 2.4.
$\operatorname{QCoh}^{Y}(X)$ and $\operatorname{QCoh}^{Y}(\mathfrak{X})$ are equivalent to $\operatorname{QCoh}(Y)$ and $\operatorname{QCoh}(\mathfrak{Y})$, respectively.
As a consequence, we get the following.
Corollary 2.5.
Let $Y$ be a closed subvariety of $X$. Then the spectral action makes $\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}_{G})$ into a module for $\operatorname{QCoh}(\mathfrak{Y})$.
For computations, we will also need to consider the $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}^{b}_{G})$. Associated with $i^{b}:\mathrm{Bun}_{G}^{b}\to\mathrm{Bun}_{G}$ we have an adjoint pair of functors $(i^{b\ast},i_{\ast}^{b})$, defined in [FS21, §VII.6]. Since $\operatorname{QCoh}(X)$ is generated by $\mathcal{O}_{X}$ under cones, shifts, retracts and filtered colimits, any idempotent complete, cocomplete stable subcategory of $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ will be preserved by the action of $\operatorname{QCoh}(X)$. We can therefore define a $\operatorname{QCoh}(X)$-action on $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b})$ by declaring that
$i^{b}_{\ast}:\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b})\to\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$
is $\operatorname{QCoh}(X)$-linear.
Proposition 2.6.
The functor $i^{b\ast}:\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})\to\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b})$ is $\operatorname{QCoh}(X)$-linear with respect to the action above. In particular, for any $Y\subseteq X$ as above, $i^{b\ast}$ and $i^{b}_{\ast}$ induce functors $\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}_{G})\to\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}^{b}_{G})$ and $\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}^{b}_{G})\to\mathcal{D}_{\mathrm{lis}}^{Y}(\mathrm{Bun}_{G})$, which we will also denote by $i^{b\ast}$ and $i^{b}_{\ast}$, respectively.
Proof.
The second part follows from the first part (and the definition, in the case of $i^{b}_{\ast}$). The first part is essentially follows from [Gai, Cor. 6.2.4], except that $\operatorname{QCoh}(X)$ is not rigid since $X$ is not quasicompact. However, the categories $\operatorname{QCoh}(X^{P})$ are rigid, so using Remark 2.2 one reduces to this case. We omit the details. ∎
For our applications, we need a criterion for objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ to have an $\mathcal{O}_{Y}$-structure, phrased in terms of Fargues–Scholze parameters. For this, we will use the following lemma.
Proposition 2.7.
Let $k$ be a field and Let $A$ and $B$ be $k$-algebras, with $A$ commutative. Let $F:\mathcal{D}(A)\to\operatorname{End}_{k}^{L}(\mathcal{D}(B))$ be a monoidal $k$-linear functor which commutes with colimits. Then:
  1. (1)
    $F$ induces a ring homomorphism $f:A\to Z(B)$, where $Z(B)$ denotes the centre of $B$, and $F$ is given by $F(M)=(N\mapsto M\otimes_{A,f}^{L}N)$, for $M\in\mathcal{D}(A)$ and $N\in\mathcal{D}(B)$.
  2. (2)
    Let $I\subseteq A$ be an ideal and let $M\in\mathcal{D}(B)$. Assume that $M$ has homology in a single degree. If the map $A\to\operatorname{End}_{D(B)}(M)$ factors through $A/I$, then $M$ has a canonical structure of an $A/I$-module, and hence lies in $\mathcal{D}(A/I\otimes_{A}^{L}B)=\operatorname{Mod}_{A/I}(\mathcal{D}(B))$.
Proof.
We start with part (1). For any $k$-algebras $R$ and $S$, the $\infty$-category of colimit-preserving functors $\mathcal{D}(R)\to\mathcal{D}(S)$ is equivalent to $\mathcal{D}(S\otimes_{k}R^{op})$, or equivalently to the derived $\infty$-category $\mathcal{D}(S,R)$ of $(S,R)$-bimodules. Here $X\in\mathcal{D}(S,R)$ corresponds to the functor $M\mapsto X\otimes_{R}^{L}M$, and for a given $F:\mathcal{D}(R)\to\mathcal{D}(S)$ the corresponding $X$ is $F(R)$, with the left $R^{op}$-module structure coming from the map
$R^{op}=\operatorname{End}_{\mathcal{D}(R)}(R)\to\operatorname{End}_{\mathcal{D}(S)}(F(R))$
induced by $F$. When $R=S$, the monoidal structure on $\operatorname{End}^{L}_{k}(\mathcal{D}(R))$ by composition corresponds to the tensor product (over $R$) on $\mathcal{D}(R,R)$. Now consider the $F$ in the statement of the proposition. Using the above, we may think of it as a monoidal functor $\mathcal{D}(A)\to\mathcal{D}(B,B)$, and hence $F(A)$ is equivalent to $B$, the unit. Applying the above remarks again, now thinking of $\mathcal{D}(B,B)$ as $\mathcal{D}(B\otimes_{k}B^{op})$, $F$ itself is given by $M\mapsto B\otimes_{A}^{L}M$, with the $A$-module structure on the $(B,B)$-bimodule $B$ being given by the homomorphism
$A=\operatorname{End}_{\mathcal{D}(A)}(A)\to\operatorname{End}_{\mathcal{D}(B,B)}(B).$
Since $A$ and $B$ are concentrated in degree 00, this map factors through the center $Z(B)=\pi_{0}(\operatorname{End}_{\mathcal{D}(B,B)}(B))$, giving the desired map $f$. Translating back from $\mathcal{D}(B,B)$ to $\operatorname{End}_{k}^{L}(\mathcal{D}(B))$ then gives the desired formula.
For part (2), by shifting we may assume that $M$ has homology in degree 00, and so lies in the heart of $\mathcal{D}(B)$ with respect to the usual t-structure. The assumption then says that $M$ has a canonical $B/IB$-module structure in the usual sense. Restriction along the canonical map $A/I\otimes^{L}_{A}B\to B/IB$ then gives a canonical $A/I\otimes_{A}^{L}B$-module structure, as desired. ∎
We then have the following corollary, which gives us elements of $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$.
Corollary 2.8.
Let $\rho$ be an irreducible admissible $G_{b}(E)$-representation (in the abelian category, not the derived category) with Fargues–Scholze parameter $\varphi$. Then $i_{\ast}^{b}\rho$ naturally lives in $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$.
Proof.
We may choose a sufficiently small compact open subgroup $K\subseteq G_{b}(E)$ such that $\rho$ is generated by its $K$-fixed vectors and we have a fully faithful exact colimit-preserving embedding $\mathcal{D}(\mathcal{H}(G,K))\subseteq\mathcal{D}(G_{b}(E))$, where $\mathcal{H}(G,K)$ is the usual Hecke algebra of bi-$K$-invariant compactly supported $\overline{\mathbb{Q}}_{\ell}$-valued functions on $G_{b}(E)$. In particular, $\mathcal{D}(\mathcal{H}(G,K))$ inside $\mathcal{D}(G_{b}(E))=\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}^{b}_{G})$ is $\operatorname{QCoh}(X)$-stable and contains $\rho$, so we may apply Proposition 2.7 to deduce that $\rho\in\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}^{b}_{G})$. By Proposition 2.6, we then have $i_{\ast}^{b}\rho\in\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$. ∎
3. Cohomology of moduli spaces of local shtukas
We now apply the constructions of the previous section to the cohomology of mouli space of local shtukas. Let $b\in B(G)$ and let $I$ be a finite set and let $\mu_{\bullet}=(\mu_{i})_{i\in I}$ be a collection of conjugacy classes of cocharacters of $G$, with reflex fields $E_{\bullet}=(E_{i})_{i\in I}$. From the introduction we have the tower $(\operatorname{Sht}_{(G,b,\mu_{\bullet}),K})_{K}$ of moduli spaces of local shtukas. The “intersection cohomology complex” of $\operatorname{Sht}_{(G,b,\mu_{\bullet}),K}$ is the object $f_{K\natural}\mathcal{S}^{\prime}_{W}$ defined before [FS21, Prop. IX.3.2], where $W\in\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is the representation with highest weight $\mu_{\bullet}=(\mu_{i})_{i\in I}$. When $\mu_{\bullet}=\mu$ is a single minuscule cocharacter this is simply the (shifted) compactly supported $\ell$-adic cohomology of the smooth rigid analytic variety $\operatorname{Sht}_{(G,b,\mu),K}$. Let $\rho$ be an irreducible admissible representation of $G_{b}(E)$, with Fargues–Scholze parameter $\varphi$. Consider the $\rho$-isotypic part
[…]
of the cohomology of the tower $(\operatorname{Sht}_{(G,b,\mu_{\bullet}),K})_{K}$. By the proof of [FS21, Prop. IX.3.2] and the argument in the proof of [HKW22, Proposition 6.4.5], we have
(1) $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong i^{1\ast}T_{V}i^{b}_{\ast}\rho,$
where $V\in\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is the dual of the representation with highest weight $\mu_{\bullet}=(\mu_{i})_{i\in I}$, and this is a bounded complex of finite length $G(E)$-representations.
Now set $\mathfrak{X}_{\varphi}:={\varphi}\times_{X}\mathfrak{X}$ (in the derived sense), viewing $\varphi$ as a closed point of $X$. To understand the cohomology, we have to analyze the expression $i^{1\ast}T_{V}i^{b}_{\ast}\rho$ more closely. By Corollary 2.8, $i^{b}_{\ast}\rho\in\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$. The operator $T_{V}$ then takes us from $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$ to $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})^{BW_{E}^{I}}$, and is given by the image of $V$ under the composition
$\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})\to\operatorname{QCoh}(\mathfrak{X}_{\varphi})^{BW_{E_{\bullet}}}\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G}))^{BW_{E_{\bullet}}},$
so to understand this better we need to understand the derived stack $\mathfrak{X}_{\varphi}$. A priori, this can be a non-classical stack, and we will have little to say about this case in what follows — to see the correct structure on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ in that case one would need a refinement of Corollary 2.8. We will focus on a particular class of parameters $\varphi$ that turn out to satisfy $\mathfrak{X}_{\varphi}\cong[\ast/S_{\varphi}]$, where $S_{\varphi}$ is the centralizer of $\varphi$, viewed as a point of $\mathfrak{X}$. In what follows, we write $\pi:X^{\Box}\to X$ for the map from $X^{\Box}$ to its GIT quotient $X$. Recall that semisimple $L$-parameters over $\overline{\mathbb{Q}}_{\ell}$ precisely correspond to closed points of $X$.
Definition 3.1.
Let $\varphi$ be a semisimple $L$-parameter over $\overline{\mathbb{Q}}_{\ell}$. Let $\pi^{-1}(\varphi)$ denote the classical (non-derived) scheme-theoretic fibre of $\pi$ at $\varphi$, and let $\pi^{-1}(\varphi)_{red}$ denote its nilreduction.
  1. (1)
    We say that $\varphi$ is strongly semisimple if $\pi^{-1}(\varphi)_{red}$ consists of a single $\widehat{G}$-orbit.
  2. (2)
    We say that $\varphi$ is generous if $\pi^{-1}(\varphi)$ is reduced and consists of a single $\widehat{G}$-orbit.
Concretely, $\varphi$ is strongly semisimple if there are no other $L$-parameters with the same semisimplification as $\varphi$. We note that strongly semisimple $\varphi$ define smooth points of $X^{\Box}$ (or, equivalently, of $\mathfrak{X}$) [^3] This follows from the description of the cotangent complex on $\mathfrak{X}$ [FS21, §VIII.2], using the duality in [FS21, Prop. VIII.2.2].. Now assume that $\varphi$ is generous. Then the underlying classical stack $\mathfrak{X}_{\varphi}^{cl}$ of $\mathfrak{X}_{\varphi}$ is $\pi^{-1}(\varphi)/\widehat{G}$, so by definition $\varphi$ is generous if and only if $\mathfrak{X}_{\varphi}^{cl}\cong[\ast/S_{\varphi}]$. In fact, we will now show that $\mathfrak{X}_{\varphi}$ is classical for generous $\varphi$.
Proposition 3.2.
Let $X^{\Box}_{ss}$ and $X_{ss}$ denote the loci of strongly semisimple $L$-parameters in $X^{\Box}$ and $X$, respectively. Then $X_{ss}$ is open, $X_{ss}^{\Box}=\pi^{-1}(X_{ss})$, and $X_{ss}^{\Box}\to X_{ss}$ is a universal geometric quotient. Moreover, if $\varphi$ is generous, then there is an open neighbourhood $W\subseteq X$ of $\varphi$ such that $\pi:\pi^{-1}(W)\to W$ is flat.
Proof.
We may prove this component by component on $X$, so let $C\subseteq X$ be a connected component and set $C^{\Box}=\pi^{-1}(C)$; this is connected since $\widehat{G}$ is. By [DHKM24, Thm. 1.8], $C$ is irreducible and reduced. If there are no strongly semisimple parameters on $C$ then the assertion is trivial, so assume that $C_{ss}:=C\cap X_{ss}\neq\emptyset$. Let $C^{\Box}=Y_{1}\cup Y_{2}\cup\dots\cup Y_{r}$ be the decomposition of $C^{\Box}$ into irreducible components. Each component is $\widehat{G}$-invariant, so the sets $\pi(Y_{i})$ are closed. Since $\pi(C^{\Box})=C$, we may without loss of generality assume that $\pi(Y_{1})=C$. Then we must have $C^{\Box}_{ss}:=X^{\Box}_{ss}\cap C^{\Box}\subseteq Y_{1}\setminus(Y_{2}\cup\dots\cup Y_{r})$, since the points in $C^{\Box}_{ss}$ are smooth.
Now set $Y=C\setminus(\pi(Y_{2})\cup\dots\cup\pi(Y_{r}))$ and $Y^{\Box}=\pi^{-1}(Y)$. These are both open and irreducible and we note that $C_{ss}^{\Box}\subseteq Y^{\Box}$ (and $C_{ss}\subseteq Y$). In particular, $Y^{\Box}$ is the GIT quotient of $Y$. Consider the set $U\subseteq Y^{\Box}$ of points whose orbits have maximal dimension; $U$ is open by [New78, Lem. 3.7(c)]. Call this maximal dimension $d^{\prime}$. We let $U^{\prime}\subseteq U$ be subset of points in $U$ whose orbits are closed; this satisfies $U^{\prime}=\pi^{-1}(W^{\prime})$ for some open $W^{\prime}\subseteq Y$ by [New78, Prop. 3.8], and $U^{\prime}\to W^{\prime}$ is a universal geometric quotient. Now consider the set $V\subseteq Y^{\Box}$ of $y\in Y^{\Box}$ for which the dimension of the local ring $\mathcal{O}_{\pi^{-1}(\pi(y)),y}$ of the fibre $\pi^{-1}(\pi(y))$ is minimal; this is an open set by [Mum99, §I.8, Thm. 3 Cor. 3]. Call this minimal dimension $d$. Let $x\in C_{ss}^{\Box}$. By definition of $d^{\prime}$, we must have $d^{\prime}\geq\dim\widehat{G}x$. On the other hand, $\dim\widehat{G}x=\dim\pi^{-1}(\pi(x))$ and $\pi^{-1}(\pi(x))$ is equidimensional by genericity, so we must have $\dim\widehat{G}x\geq d$ by the definition of $d$. In particular, we have $d\leq d^{\prime}$. On the other hand, we have $U\cap V\neq\emptyset$ since $U$ and $V$ are open and $Y^{\Box}$ is irreducible, so choose a point $y\in U\cap V$. Then $d=\dim\mathcal{O}_{\pi^{-1}(\pi(y)),y}$ and $d^{\prime}=\dim\widehat{G}y$ by construction. Since $y\in\widehat{G}y\subseteq\pi^{-1}(\pi(y))$, we must have $d^{\prime}\leq d$, and hence we conclude that $d=d^{\prime}$ and that $C_{ss}\subseteq U$. Since the orbits of semisimple $L$-parameters are closed, we conclude that $C_{ss}^{\Box}\subseteq U^{\prime}$. The converse also holds: Any orbit in $U^{\prime}$ is closed and cannot be in the closure of any other orbit (since such an orbit would need to have dimension bigger than $d^{\prime}$ and lie in $Y^{\Box}$, which is impossible), so we conclude that $C_{ss}^{\Box}=U^{\prime}$ and hence that $W^{\prime}=C_{ss}$. In particular $C_{ss}$ and $C_{ss}^{\Box}$ are open, and $C_{ss}^{\Box}\to C_{ss}$ is a universal geometric quotient. This finishes the proof of the first part of the proposition.
For the second part, note that we have shown that $C_{ss}^{\Box}\to C_{ss}$ is equidimensional (with fibres of dimension $d$) and universally open (since it is a universal geometric quotient), and we know that $C_{ss}$ is reduced since $C$ is. So assume that $\varphi$ is a generous parameter in $C$. Then $\pi^{-1}(\varphi)$ is reduced by assumption, so by [Gro66, Cor. (15.2.3)], it follows that $C_{ss}^{\Box}\to C_{ss}$ is flat in a neighbourhood of $\pi^{-1}(\varphi)$. We may then spread out this neighbourhood using the $\widehat{G}$-action (since we have a geometric quotient) and take the image under $\pi$ to obtain the desired $W$. ∎
Corollary 3.3.
If $\varphi$ is a generous $L$-parameter, then $\mathfrak{X}_{\varphi}=\mathfrak{X}_{\varphi}^{cl}=[\ast/\widehat{G}]$. Moreover, for any strongly semisimple $L$-parameter $\varphi$, $\pi$ is flat in a neighbourhood of $\varphi\in X$ if and only if $\varphi\in X$ is smooth.
Proof.
We have already proved the first part. For the second part, we may work component by component, and we use the notation from the proof of Proposition 3.2. There we have showed that $\pi:C^{\Box}_{ss}\to C_{ss}$ is equidimensional, and we know that $C^{\Box}_{ss}$ is smooth. That smoothness of $C_{ss}$ at $\varphi$ implies flatness of $\pi$ at $\varphi\in C_{ss}^{\Box}$ then follows from miracle flatness [Mat89, Thm. 23.1], and the converse follows from [Mat89, Thm. 23.7(i)]. ∎
Remark 3.4.
All strongly semisimple $L$-parameters for $\operatorname{GL}_{n}$ are generous. On the other hand, let $\zeta$ be a primitive $n$-th root of unity and consider the unramified $L$-parameter $\varphi_{n}$ sending Frobenius to $(1,\zeta,\dots,\zeta^{n-1})\in\operatorname{PGL}_{n}(\overline{\mathbb{Q}}_{\ell})$. Then $\varphi_{n}$ is strongly semisimple but not generous, and it is a smooth point in $X$ if and only if $n=2$. We also remark that there is a Zariski open and dense subset of generous parameters inside $X_{ss}$ by [Sta24, Tag 0578] (the full set of generous parameters is locally constructible by [Sta24, Tag 0579], but we do not know if it is open).
We now return to analyzing $R\Gamma(G,b,\mu_{\bullet})[\rho]$, under the assumption that the Fargues–Scholze parameter $\varphi$ of $\rho$ is generous. Since $\mathfrak{X}_{\varphi}\cong[\ast/S_{\varphi}]$, we have $\operatorname{QCoh}(\mathfrak{X}_{\varphi})\cong\operatorname{Rep}(S_{\varphi})$ and the action of $V$ on $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$ factors as
(2) $\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})\to\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})^{\omega})^{BW_{E_{\bullet}}}.$
By the compatibility between $b$ and $\mu_{\bullet}$, $Z(\widehat{G})^{Q}\subseteq S_{\varphi}$ acts on $V$ via a character $\chi_{b}$. Write $\operatorname{Irr}(S_{\varphi},\chi_{b})$ for the set of irreducible representations of $S_{\varphi}$ on which $Z(\widehat{G})^{Q}$ acts by $\chi_{b}$; this is a finite set. Also write $V_{\varphi}$ for the underlying $W_{E_{\bullet}}$-representation on the image of $V$ under $\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})\to\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}$. We can now put everything together to derive the desired formula.
Theorem 3.5.
Assume that the Fargues–Scholze parameter $\varphi$ of $\rho$ is generous, and decompose the image of $V$ in $\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}$ as $\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}\delta\boxtimes V_{\varphi,\delta}$. Then we have an isomorphism
$R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)\boxtimes V_{\varphi,\delta}$
in $\mathcal{D}^{\varphi}(G(E))^{BW_{E_{\bullet}}}$, and each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a bounded complex of finite length $G(E)$-representations.
Proof.
The formula follows immediately from equations (1) and (2), and the last statement follows from the fact that $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a direct summand of $R\Gamma(G,b,\mu_{\bullet})[\rho]$. ∎
Remark 3.6.
We make some remarks on Theorem 3.5.
  1. (1)
    We expect that the theorem holds when $\varphi$ is strongly semisimple. For this, one would want to show that $i^{b}_{\ast}\rho$ has a $\operatorname{QCoh}([\ast/S_{\varphi}])$-structure, but this is stronger than $\rho$ having Fargues–Scholze parameter $\varphi$ when $\varphi$ is not generous. In situations when $\varphi$ lifts to a generous parameter on an isogenous group (as in e.g. Remark 3.4), we expect that one can prove this stronger statement, but we have not checked the details.
  2. (2)
    The categorical conjecture predicts that $\mathcal{D}^{\varphi}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ is equivalent to $\mathcal{D}(\operatorname{Rep}(S_{\varphi}))$ when $\varphi$ is generous. In particular, each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ should be a split complex.
  3. (3)
    Each $\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is non-zero but, unless $\varphi$ is elliptic, many $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ will be zero (indeed, only finitely many can be non-zero).
  4. (4)
    With appropriate modifications, our basic definitions go through with integral coefficients, but making the arguments go through would require more work.
  5. (5)
    Let us briefly compare our results to those of Koshikawa [Kos21], in particular Thm. 1.3 of loc. cit, which says that the $W_{E_{\bullet}}$-representations appearing in $R\Gamma(G,b,\mu_{\bullet})[\rho]$ are subquotients of $V_{\varphi}$, without any assumption on $\varphi$ and allowing integral coefficents. In characteristic 00, we can recover this statement from the assertion that $i^{1\ast}T_{V}i_{b}^{\ast}\rho\in\mathcal{D}^{\varphi}(G(E))^{BW_{E\bullet}}$ (which does not require any assumption on $\varphi$). Moreover, a formula of the form $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}C_{\delta}\boxtimes V_{\varphi,\delta}$ can be deduced straight from the spectral action as in the proof of [Kos21, Thm. 1.3], without the machinery developed here. Indeed, the spectral action shows that the $W_{E_{\bullet}}$-action on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ factors as
    ¯$\overline{\mathbb{Q}}_{\ell}[W_{E_{\bullet}}]\to\operatorname{End}(\mathcal{V})/\mathfrak{m}_{\varphi}\to\operatorname{End}(R\Gamma(G,b,\mu_{\bullet})[\rho]),$
    where $\mathfrak{m}_{\varphi}\subseteq\mathcal{Z}^{\mathrm{spec}}(G)$ is the maximal ideal cutting out $\varphi$ and $\mathcal{V}$ is the $W_{E_{\bullet}}$-equivariant vector bundle on $\mathfrak{X}$ corresponding to $V$. When $\varphi$ is generous, one has $\operatorname{End}(\mathcal{V})/\mathfrak{m}_{\varphi}=\operatorname{End}_{S_{\varphi}}(V_{\varphi})$, and by standard representation theory one gets the desired decomposition. The extra information gained from factoring the spectral action is the functorial formula for the $C_{\delta}$.
In the rest of this section, we address points (2) and (3) of Remark 3.6 when $\varphi$ is elliptic. Recall that $\varphi$ is said to be elliptic if it is semisimple and $S_{\varphi}/Z(\widehat{G})^{Q}$ is finite. In this case, the discussion in [FS21, §X.2] shows that, for any $b\in B(G)$, a $G_{b}(E)$-representation $\sigma$ with Fargues–Scholze parameter $\varphi$ has to be supercuspidal, and $b$ has to be basic. In particular,
(3) $\prod_{b\in B(G)_{\mathrm{basic}}}i^{b\ast}:\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})\longrightarrow\prod_{b\in B(G)_{\mathrm{basic}}}\mathcal{D}^{\varphi}(G_{b}(E))$
is an equivalence. We then have the following assertion.
Lemma 3.7.
Assume that $\varphi$ is elliptic and that $b$ is basic. Then $\mathcal{D}^{\varphi}(G_{b}(E))$ is a product of categories of the form $\mathcal{D}(A)$, for semisimple (not necessarily commutative) Artinian rings $A$. In particular, any compact object in $\mathcal{D}^{\varphi}(G_{b}(E))$ is equivalent to a finite direct sum of shifted supercuspidal representations. Moreover, $\varphi$ is generous.
Proof.
Let $C\subseteq X$ be the connected component containing $\varphi$, let $C^{\Box}$ be its preimage on $X^{\Box}$, and let $Y_{ur}$ the group variety of unramified characters $W_{E}\to Z(\widehat{G})\rtimes Q$. By the local Langlands correspondence for tori, $Y_{ur}$ is isomorphic to the group variety of unramified smooth characters of $G_{b}(E)$. Moreover, $C$ consists of all twists of $\varphi$ by elements of $Y_{ur}$, and all these are generic. In particular, $C^{\Box}\to C$ is smooth with fibres isomorphic to $\widehat{G}/S_{\varphi}$, showing that $\varphi$ is generous.
Now consider $\mathcal{D}^{C}_{\mathrm{lis}}(\mathrm{Bun}_{G})$, which is equivalent to the product of the $\mathcal{D}^{C}(G_{b}(E))$ for $b\in B(G)_{\mathrm{basic}}$. Each $\mathcal{D}^{C}(G_{b}(E))$ is the product of its Bernstein components, all of which are supercuspidal, and the action of $\operatorname{QCoh}(C)$ preserves the Bernstein components. Let $\mathcal{C}$ be such a Bernstein component, let $\mathcal{Z}(\mathcal{C})$ denote its center and let $\mathcal{R}(\mathcal{C})$ denote the endomorphism ring of a compact generator of $\mathcal{C}$. We have that $\operatorname{QCoh}(C)$ is equivalent to $\mathcal{D}(\mathcal{O}(C))$ and $\mathcal{C}$ is equivalent to $\mathcal{D}(\mathcal{R}(\mathcal{C}))$. Thus, by Proposition 2.7, the action of $\operatorname{QCoh}(C)$ on $\mathcal{C}$, viewed as an action of $\mathcal{D}(\mathcal{O}(C))$ on $\mathcal{D}(\mathcal{R}(\mathcal{C}))$, is given by letting $M\in\mathcal{D}(\mathcal{O}(C))$ act by the endomorphism
$N\mapsto M\otimes_{\mathcal{O}(C),f}N$
on $\mathcal{D}(\mathcal{R}(\mathcal{C}))$, where $f:\mathcal{O}(C)\to\mathcal{Z}(\mathcal{C})$ is the induced homomorphism on centres. Both $\mathcal{O}(\mathcal{C})$ and $\mathcal{Z}(\mathcal{C})$ carry twisting actions of $Y_{ur}$; indeed by choosing a base point they are both isomorphic to quotients of $Y_{ur}$ by finite groups. Since since the Fargues–Scholze construction is compatible with twisting [FS21, Thm. IX.0.5(ii)], $f$ is equivariant for the actions of $Y_{ur}$, and hence finite Galois.
Armed with this, we now consider $\mathcal{D}^{\varphi}(G_{b}(E))$, which is the product of the categories $\mathcal{C}^{\varphi}:=\operatorname{Mod}_{\mathcal{O}(C)/\mathfrak{m}_{\varphi}}\mathcal{C}$, where $\mathcal{C}$ ranges over the Bernstein components of $\mathcal{D}^{C}(G_{b}(E))$ and $\mathfrak{m}_{\varphi}\subseteq\mathcal{O}(C)$ is the maximal ideal corresponding to $\varphi$. By the description of the action of $\operatorname{QCoh}(C)$ on $\mathcal{C}$, $\mathcal{C}^{\varphi}$ is equivalent to $\mathcal{D}(\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C}))$, so we need to show that $\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C})$ is a (classical) semisimple Artinian ring. By [BH03, 8.1, Prop.], $\mathcal{R}(\mathcal{C})$ is an Azumaya algebra over $\mathcal{Z}(\mathcal{C})$. Since $f$ is finite Galois, it follows that $\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C})$ is concentrated in degree 00, where it is a finite product of matrix algebras over $\overline{\mathbb{Q}}_{\ell}$, and in particular semisimple and Artinian, as desired. The rest of the lemma then follows immediately. ∎
We now get the following refinement of Theorem 3.5 for elliptic parameters.
Corollary 3.8.
In the situation of Theorem 3.5, assume that $\varphi$ is elliptic. Then each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a non-zero finite direct sum of shifted supercuspidal representations. If $\delta$ is one-dimensional, then $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a single supercuspidal representation, up to shift.
Proof.
$\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is non-zero and supported on $\mathrm{Bun}_{G}^{1}$ by equation (3), hence $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is non-zero. It is a finite direct sum of supercuspidal representations up to shift by Lemma 3.7. Finally, if $\delta$ is one-dimensional, then $\operatorname{Act}_{\delta}$ is an equivalence, so $\operatorname{End}(i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho))=\operatorname{End}(\rho)=\overline{\mathbb{Q}}_{\ell}$ by the previous observation on the support. It follows that $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a single supercuspidal representation up to shift. ∎
Remark 3.9.
We note that Corollary 3.8 has previously appeared in the literature; see [FS21, §X.2] for the case when $S_{\varphi}$ is finite and [Ham22, Cor. 3.11], [BMHN22, Thm. 2.27] for the general case, where it plays a key role in comparing the Fargues–Scholze construction with the local Langlands correspondence. Note that, in light of Corollary 3.8, the Kottwitz conjecture (with respect to the Fargues–Scholze construction) amounts to the image of $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ in the Grothendieck group of $G(E)$ being the supercuspidal representation in the $L$-packet of $\varphi$ corresponding to $\delta$, in the parametrization that uses $\rho$ as a base point. The vanishing conjecture, on the other hand, says that $i_{1}^{\ast}\operatorname{Act}_{\delta}(i_{b\ast}\rho)$ is concentrated in degree 00. In particular, Corollary 3.8 gives a complete understanding of the action of the Weil group. This is in contrast to local approaches to the cohomology of local Shimura varieties predating [FS21], which could say very little about the Weil group action.
4. Some consequences
In this section we indicate some results towards Conjecture 1.1 that can be obtained from Theorem 3.5 using simple tricks or observations, or recent works on Kottwitz conjecture. We keep the notation and assumptions from §3, and assume additionally that $\varphi$ is elliptic. In this case, the set $\operatorname{Irr}(S_{\varphi},\chi_{b})=\{\delta_{1},\dots,\delta_{r}\}$ is finite, and we write $V_{\varphi,i}$ for $V_{\varphi,\delta_{i}}$. We also assume that $E$ has characteristic 00 throughout this section.
4.1. The Kottwitz conjecture
Let us start with the simplest possible version of Corollary 3.8: Assume that $\varphi$ is stable, i.e. $S_{\varphi}=Z(\widehat{G})^{Q}$. Then the formula reads
$R\Gamma(G,b,\mu_{\bullet})[\rho]\cong i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)\boxtimes V_{\varphi},$
and $i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)$ is a single supercuspidal representation up to shift. The following proposition is then clear (note that $\chi_{1}$ is the trivial representation).
Proposition 4.1.
Assume that $\varphi$ is stable. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]$ vanishes outside a single degree, and in that degree it is the exterior tensor product of a supercuspidal representation and $V_{\varphi}$. If, additionally, $b=1$, then $R\Gamma(G,b,\mu_{\bullet})[\rho]=\rho\boxtimes V_{\varphi}$, i.e. the strong Kottwitz conjecture holds.
For the rest of this section we proceed with a weaker assumption on $\varphi$: We assume that all the $\delta_{i}$ are one-dimensional. This is satisfied in many examples. As noted in the introduction, the local Langlands conjecture predicts the existence of an $L$-parameter $\phi$ attached to $\rho$, and an $L$-packet $\Pi_{\phi}(G)$ of $G(E)$-representations attached to $\phi$. Note that $\Pi_{\phi}(G)$ is a multiset. Write
$i^{1\ast}\operatorname{Act}_{\delta_{i}}(i^{b}_{\ast}\rho)=\pi_{i}^{\prime}[n_{i}]$
where $\pi_{i}^{\prime}$ is supercuspidal and $-[n_{i}]$ denotes a shift by an integer $n_{i}$. In [HKW22], the Kottwitz conjecture is proven modulo the action of $W_{E_{\bullet}}$, under further assumptions on $L$-packets. Combining this with Theorem 3.5, we get Corollary 1.3 from the introduction.
Theorem 4.2.
Assume the refined Local Langlands conjecture [Kal16, Conjecture G] and let $\phi$ be the $L$-parameter attached to $\rho$. Assume further that the $L$-packet $\Pi_{\phi}(G)$ only consists of supercuspidal representations, and that all $\delta_{i}$ are one-dimensional. Then
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{i=1}^{r}\pi_{1}^{\prime}\boxtimes V_{\varphi,i},$
and $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\phi}(G)$ as sets. Moreover, $H^{n}(R\Gamma(G,b,\mu_{\bullet})[\rho])=0$ when $n$ is odd.
Proof.
By Theorem 3.5, $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{i=1}^{r}(-1)^{n_{i}}\pi_{i}^{\prime}\boxtimes V_{\varphi,i}$ in $\mathrm{Groth}(G(E)\times W_{E_{\bullet}})$. On the other hand, $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{j=1}^{s}\pi_{j}\cdot\dim V_{\phi,j}$ in $\mathrm{Groth}(G(E))$ by [HKW22, Thm. 1.0.2], where $\{\pi_{1},\dots,\pi_{s}\}=\Pi_{\phi}(G)$ as multisets. By comparing these expressions we see that the $n_{i}$ are even and that $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\varphi}(G)$ as sets. In this comparison, we have used that $\sum_{j=1}^{s}\dim V_{\phi,i}=\dim V=\sum_{i=1}^{r}\dim V_{\varphi,i}$ and that, for any $j$, there exists a $\mu_{\bullet}$ such that $V_{\phi,j}\neq 0$, and similarly for $V_{\varphi,i}$. This proves the theorem. ∎
Remark 4.3.
We make a few remarks on this theorem.
  1. (1)
    While [HKW22, Thm. 1.0.2] is only stated for $\mu_{\bullet}=\mu$ a single cocharacter, its proof goes through for general $\mu_{\bullet}$. We note, however, that Theorem 4.2 for general $\mu_{\bullet}$ follows from the case $\mu_{\bullet}=\mu$, as this case suffices to show that $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\phi}(G)$, which is the content of Theorem 4.2 in view of Theorem 3.8.
  2. (2)
    As noted in the introduction, this is rather close to the Kottwitz conjecture, but falls short in two aspects. The first issue is that we do not know that $\phi=\varphi$ (indeed, it is currently not known that $\varphi$ is elliptic in most cases when we expect it to be). Lacking this our result looks rather amusing: The $G(E)$-representations in $R\Gamma(G,b,\mu_{\bullet})[\rho]$ arise from $\phi$, whereas the $W_{E_{\bullet}}$-action is given in terms of $\varphi$. The second issue is that, even if $\phi=\varphi$, it is not clear to us that the natural parametrizations of $\Pi_{\phi}(G)$ are the same. At present, the only approach available to these questions is through the cohomology of (global) Shimura varieties. It is known that $\phi=\varphi$ for inner forms of $\operatorname{GL}_{n}$ and $\operatorname{SL}_{n}$ for general $E$ [FS21, HKW22], for inner forms of $\operatorname{GSp}_{4}$ and $\operatorname{Sp}_{4}$ when $E$ is unramified over $\mathbb{Q}_{p}$ [Ham22], for some (similitude) unitary groups in an odd number of variables [BMHN22] when $E=\mathbb{Q}_{p}$, and for $\operatorname{SO}_{2n+1}$ and its unique inner form when $E$ is unramified over $\mathbb{Q}_{p}$ (D.H., unpublished). Comparing the parametrizations of the $L$-packet seems to be more subtle, however.
  3. (3)
    We also remark that the extra assumptions of Theorem 4.2 (excluding that $\varphi$ is elliptic) are known to hold in many cases. Examples include inner forms of $\operatorname{SL}_{n}$, $\operatorname{Sp}_{2n}$, $\operatorname{SO}_{2n+1}$ and unitary groups. Moreover, [Kal16, Conjecture G] is known for the regular supercuspidal $L$-packets constructed by Kaletha for all $G$ split over a tame extension of $E$ and for $p$ sufficiently large, in [Kal19]; see [Kal19, FKS23]. We refer to the introduction of [HKW22] for more details.
4.2. Vanishing
One interesting aspect of Theorem 3.5 is its uniformity when varying $\mu_{\bullet}$: Only the $W_{E_{\bullet}}$-representations change [^4] A trivial consequence is that, for fixed $(G,b)$ and $\rho$, there exists an $N$ such that $H^{i}(R\Gamma(G,b,\mu_{\bullet})[\rho])$ vanishes for all $|i|>N$ and all $\mu_{\bullet}$.. This observation can sometimes be used to propagate vanishing results for some $\mu_{\bullet}$ to a larger collection (Conjecture 1.1(2) was recently proven by the first author in [Han21] in many cases where $\mu_{\bullet}$ is a single minusucle cocharacter). To illustrate the method, we reprove the strong Kottwitz conjecture for inner forms of $\operatorname{GL}_{n}$, which was previously proved by the first author [Han21, Thm. 1.9], using results on averaging functors from [ALB21].
Theorem 4.4.
Assume that $G$ is an inner form of $\operatorname{GL}_{n}$. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\pi\boxtimes V_{\varphi}$, where the right hand is concentrated in degree 00 and $\pi$ is the irreducible admissible representation of $G(E)$ corresponding to $\varphi$ under the usual local Langlands correspondence.
Proof.
Note that we are in the stable situation here, so we need to show that $i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)$ is $\pi$ concentrated in degree 00. That it is a shift of $\pi$ follows from [HKW22, Thm. 1.0.3], since its Fargues–Scholze $L$-parameter is $\varphi$. To prove vanishing, note that we can choose a minuscule cocharacter $\mu_{b}$ such that $R\Gamma(G,b,\mu_{b})[\rho]=i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)\boxtimes W_{\varphi}$, where $W$ the dual of the irreducible representation with extreme weight $\mu_{b}$. By [Han21, Thm. 1.6], $R\Gamma(G,b,\mu_{b})[\rho]$ vanishes outside degree 00, finishing the proof. ∎
Another trick that can sometimes be used is duality. Letting $\mathbb{D}$ denote Verdier duality on $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$, one has
$\mathbb{D}(R\Gamma(G,b,\mu_{\bullet})[\rho])\cong\mathbb{D}(i^{1\ast}T_{V}i^{b}_{\ast}\rho)\cong i^{1\ast}T_{V}i^{b}_{\ast}\rho^{\vee}\cong R\Gamma(G,b,\mu_{\bullet})[\rho^{\vee}],$
as objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ (i.e. forgetting the $W_{E_{\bullet}}$-action), where we have used that $i_{1}^{\ast}=i_{1}^{!}$, $i_{b!}\rho^{\vee}=i_{b\ast}\rho^{\vee}$, and the interplay between $\mathbb{D}$ and pullback/pushforward, and the Hecke operators [FS21, Thm. IX.0.1(i)]. This gives us the following vanishing theorem. In its formulation, we note that there is a canonical isomorphism between the cocenters of $G$ and $G_{b}$ over $E$, so any smooth character of $G(E)$ can be naturally viewed as a smooth character of $G_{b}(E)$ (and vice versa).
Proposition 4.5.
Assume that the $\delta_{i}$ are one-dimensional and that, writing $i^{1\ast}\operatorname{Act}_{\delta_{i}}(i^{b}_{\ast}\rho)=\pi_{i}[n_{i}]$, the $\pi_{i}$ are distinct. Assume further that there is a smooth character $\chi$ of $G(E)$ such that $\rho^{\vee}\cong\rho\otimes\chi$ and $\pi_{i}^{\vee}\cong\pi_{i}\otimes\chi$ for all $i$. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]$ is concentrated in degree 00.
Proof.
Since $\rho^{\vee}\cong\rho\otimes\chi$, Verdier duality gives us that $\mathbb{D}(R\Gamma(G,b,\mu_{\bullet})[\rho])\cong R\Gamma(G,b,\mu_{\bullet})[\rho\otimes\chi]$. Computing from the definitions, one sees that $R\Gamma(G,b,\mu_{\bullet})[\rho\otimes\chi]\cong R\Gamma(G,b,\mu_{\bullet})[\rho]\otimes\chi$. Thus, evaluating both sides using Theorem 3.8, we get that
$\bigoplus_{i=1}^{r}\pi_{i}^{\vee}[-n_{i}]^{\oplus\dim V_{\varphi,i}}\cong\bigoplus_{i=1}^{r}(\pi_{i}\otimes\chi)[n_{i}]^{\oplus\dim V_{\varphi,i}}.$
Since $\pi^{\vee}_{i}\cong\pi_{i}\otimes\chi$ and the $\pi_{i}$ are assumed to be distinct, it follows that $n_{i}=0$ for all $i$ as desired. ∎
Remark 4.6.
Let us finish with a few remarks on Proposition 4.5, and the vanishing conjecture more generally.
  1. (1)
    In the stable case, the assumption $\pi_{i}^{\vee}\cong\pi_{i}\otimes\chi$ can be dropped, and instead deduced as a consequence of the argument.
  2. (2)
    That the $\pi_{i}$ are distinct follows from the refined local Langlands correspondence whenever $G$ is an extended pure inner form of its quasi-split inner form. It may fail when this is not the case, for example when $G$ is the non-split inner form of $\operatorname{SL}_{2}$ and $S_{\varphi}=(\mathbb{Z}/2)^{2}$, in which case the $\pi_{i}$ are all equal.
  3. (3)
    Essential self-duality is a fairly common feature among representations of classical groups and related groups. For example, it holds for all representations of $\operatorname{GSp}_{2g}$; see [Pra19, Remark 5] (in this case $\chi$ is determined by the central character and hence by $\varphi$). Other examples include $\operatorname{SO}_{2n+1}$ and the unique non-split inner form of $\operatorname{GSp}_{4}$.
  4. (4)
    Conjecture 1.3(2) can sometimes be passed through isogenies. For example, one can deduce it for $\operatorname{SL}_{n}$ and $\operatorname{Sp}_{4}$ from the case of $\operatorname{GL}_{n}$ and $\operatorname{GSp}_{4}$, respectively, using computations as in the proof of [FS21, Thm. IX.6.1].